Under certain circumstances, potassium ions (K+) in a cell will move across the cell membrane from the inside to the outside. The potential inside the cell is-80.5 mV and the potential outside the cell is zero. What is the change in the electrical potential energy of a single potassium ion as it moves across the membrane? change in potential energy: Does the potential energy of the potassium ions increase, decrease, or remain the same as the ions move from the inside to the outside? D increase O remain the same O decrease
Answer #1

DU = 9x DV = 1.6810-19% (0-(-80.5*10-3)) = 1.288 × 10-20 J Ausun Potential Energy increase